The cost of building an office complex, x floors high, in a prime location in Accra is made up of three components:


GH¢10 million for the land

GH¢1/4 million per floor

Specialized costs of GH¢10000x per floor.


How many floors should the office complex contain if the average cost per floor is to be minimized?


1
Expert's answer
2021-05-20T11:35:00-0400

Average cost per floor "ATC" is


"ATC=\dfrac{TC}{x}=\dfrac{10}{x}+0.25+0.01x ,"

million per floor


Marginal cost is


"MC=(10+0.25x+0.01x^2)'=0.25+0.02x"

million per floor

Average cost per floor "ATC" is minimized if "ATC=MC"


"\dfrac{10}{x}+0.25+0.01x =0.25+0.02x,\ x>0"

"\dfrac{10}{x}-0.01x =0"

"1000-x^2=0"

"x=\sqrt{1000}"

"x=31"


"ATC=\dfrac{10}{31}+0.25+0.01(31)=0.8826"

"x=32"



"ATC=\dfrac{10}{32}+0.25+0.01(32)=0.8825"

The office complex should contain 32 floors if the average cost per floor is to be minimized.